8) Calculate the following for 2.0 L solution containing [Ag+] = 0.100 M and [Pb2+] = 0.100 M. Assume no volume changes. (AgCl Ksp = 1.8 x 10-10, PbCl2 Ksp = 1.8 x 10-5).

a) At what [Cl-] will each salt precipitate?

b) What percent of the Ag+ has precipitated before the Pb2+ begins to precipitate?

c) How much sodium chloride must be added (in grams) to precipitate a maximum AgCl before before any PbCl2 begins to precipitate?
Answer #1

Answer – We are given, [Ag+] = 0.100 M and [Pb2+] = 0.100 M

AgCl Ksp = 1.8 x 10-10, PbCl2 Ksp = 1.8 x 10-5

a) For AgCl –

We know Ksp expression for AgCl

Ksp = [Ag+] [Cl-]

1.8 x 10-10 = 0.100 M *[Cl-]

[Cl-] = 1.8 x 10-10 / 0.100 M

        = 1.8*10-9 M

so, we need 1.8*10-9 M [Cl-] to start AgCl precipitate.

For PbCl2

We know Ksp expression for PbCl2

Ksp = [Pb2+] [Cl-]2

1.8 x 10-5 = 0.100 M *[Cl-]

[Cl-]2 = 1.8 x 10-5 / 0.100 M

        = 1.8*10-4 M

squar root from both side

[Cl-] = 0.0134 M

so we need 0.0134 M [Cl-] to start PbCl2 precipitate.

b) For the bigan precipitation of PbCl2 there is needed 0.0134 M Cl-

we need to calculate for the 0.0134 M Cl- how much Ag+ remaining

Ksp = [Ag+] [Cl-]

1.8 x 10-10 = [Ag+] * 0.0134 M

[Ag+] = 1.8 x 10-10 / 0.0134

           = 1.34 *10-8 M

so there is remaining in solution [Ag+] =1.34 *10-8 M

Ag+ used in precipitation = 0.100 - 1.34 *10-8 M = 0.100 M

Means almost all 100 % Ag+ is precipitate before the Pb2+ begins to precipitate.

c) We know for the start precipitation of PbCl2 we need more than concentration of 0.0134 M Cl-

Means at 0.0134 M Cl- there are all Ag+ precipitate, since only 1.8 *10-9 M Cl- required to start precipitation

so, [Cl-] =[NaCl] = 0.0134 M

Moles of NaCl = 0.0134 M x 2.0 L

                        = 0.0268 moles

Mass of NaCl = 0.0268 moles * 58.44 g/mol

                       = 1.57 g

so, 1.57 g of sodium chloride must be added (in grams) to precipitate a maximum AgCl before any PbCl2 begins to precipitate.